# The temperature and metallicity distributions of the ICM: insights with TNG-Cluster for XRISM-like observations

### v1

@@ 4.1.2 Dependence of quality of the fit on number of photons
In Figure 9 we examine this hypothesis by comparing each model inference on the temperature and Fe abundance as a function of the number of counts. It is evident from the distribution of our clusters that the rejected bvlognorm best-fits tend to have systematically lower counts than the accepted ones, but the count distributions of the two populations have a significant overlap. We note that the same behaviour holds for the accepted and rejected best-fits of the other spectral emission models.
@@
To understand the extent that exposure time may have on the rejection of our best-fits, we repeat the X-ray spectra mocking and analysis for the rejected clusters by increasing the exposure time to 300 ks, instead of the fiducial 100 ks adopted throughout. For the case of the bvlognrom model (which returns the largest success rate), we find that 54 per cent of the clusters that have been previously rejected turn out to have a valid fit with an increased exposure (not shown). This confirms the hypothesis that the rejection of a spectral fit for the fainter clusters in our sample is due to available exposure time.
@@
However, this is not just a cluster-mass effect, as in fact there still remains a population of systems for which the best-fit is still rejected, even though their inferences on the emission-weighted mean temperature and Fe abundance are comparable with the rest of the sample. We find (although do not show) that clusters that are rejected even for a three-time longer exposure are metal poor, with the majority of the emission-weighted mass in the ICM having an Fe abundance of 0.25 Solar – such average spans between 0.15 and 0.9 Solar across the whole TNG-Cluster systems, e.g. Figure 7, bottom left panel.
@@
This is probably due to the fact that the spectral-emission models use the emission lines also to recover the thermal distribution of the gas. Whereas the shape of the continuum is a reliable diagnostic of the ICM’s bulk temperature, for plasma with temperatures higher than 2 keV, the emission-line ratios trace the presence of cooler or hotter gas, constraining the width of the distribution. However, if clusters have low elemental abundances, they will return low numbers of counts in spectral lines irrespective of the longer exposures, and in turn the modelling will be unable to provide any meaningful constraints on the width of the thermal distribution, leading to their rejection.
@@
Overall this indicates that even the more accurate assumption of a Log-Normal temperature distribution is not a sufficient representation of the properties of the ICM, as predicted by TNG-Cluster, both for its thermal and more importantly its chemical structure. This highlights the limitations of current spectral emission models and the need for more sophisticated descriptions of the intrinsic properties of the gas in order to better understand the ICM.
@@ 4.2 Effects of observational realism
Although it may be assumed in observations-based works that the observed spectra are representative of the de-projected, local conditions of the ICM that a telescope is pointing at, our analysis above suggests that there are systematic biases between the true mass-weighted state of the ICM that truly resides at the centre of clusters and the emission-weighted summary statistics of all the ICM gas that contributes to the signal from a central-pointing observation: see differences in the inference errors between top and bottom panels of e.g. Figure 5 and Figure 7. We now explore the individual sources of those differences and quantify their impact and systematic uncertainties. For the purposes of this analysis we focus again exclusively on the bvlognorm model as a reference point. However, the qualitative effects that we find are consistent across all models.
@@ 4.2.1 Projection effects
In observations, all gas in projection along the line-of-sight (LoS) contributes to the observed spectra, even though it may not be associated with the central ICM that we are interested in here. Previous works using simulated clusters have shown that projection effects can lead to severe misinterpretations of the intrinsic thermal and chemical properties of the ICM due to the gas along the LoS (e.g. Zhang et al., 2024; Štofanová et al., 2024). In order to quantify the impact of the projected gas in our work we compare the mass-weighted distributions of all the cluster gas in a central pointing with a XRISM-like FoV to the mass-weighted distribution of the gas cells that actually reside in the region of interest, in this case a sphere of a 33.5 kpc radius located in the centre of each cluster.
@@
In Figure 10, we compare the thermal structure of the gas in the centre to the total gas observed in our mocks (top right), and their corresponding inference errors (top left). Since the temperature distribution of the gas in the cluster centre can be approximated by a unimodal distribution, we bin the clusters in percentiles based on the mode of their temperature distribution. In fact, due to the tight correlation between ICM temperature and the halo mass, the temperature bins of Figure 10 top right are comparable to mass bins. Differently than e.g. in Figure 5, here we isolate the effects of projections alone, by comparing directly mass-weighted vs. mass-weighted quantities rather than mass-weighted vs. emission weighted ones.
@@
We see that the projection of gas from larger radii along the LoS shifts the temperature distribution towards lower temperatures (dashed vs. solid curves), due the projection of cooler gas from the outskirts of the cluster. The effect is more pronounced for systems with higher ICM temperature, or equivalently higher halo mass, due to the stronger temperature gradient up to in those systems. This is consistent with the strong dependence of the temperature inference on halo mass, found in Figure 5. By comparing the difference between the inferred temperature with the bvlognorm model on the one side and the the mean, mass-weighted temperature of the central vs. LoS gas, we find much larger inference errors in the former case, with average offsets of -1.2 keV (-18 per cent) and 0.16 keV (3 per cent), respectively, with a strong dependence on halo mass. Given that, on average, the inferred ICM effective temperature differs by -1.2 keV from the true one (Figure 5, and Table 2), we can conclude that projection effects dominate the observational bias.
Due to the projection of cooler gas, the width of the distribution systematically increases for all systems, but the effect shows a similar positive correlation with halo mass, being more pronounced for the most massive systems. While the increase is asymmetric, i.e. primarily for cooler gas, the tails of sub-keV gas remain unaffected overall, highlighting their central origin. However, with or without projected gas, the capability of the bvlognorm model to accurately return the true width of the input distribution remain unchanged (not shown).
@@
Projection effects are far more influential on the chemical structure of the ICM. In Figure 10 (bottom) we show the mass-weighted Fe abundance distributions of both central and projected gas. Similar to the temperature distributions, we bin the clusters based on the mode of the central gas, as it consists of one prominent peak. However, unlike the temperature of the ICM, TNG-Cluster shows no clear correlation between halo mass and metal enrichment. Our results indicate that, due to projection effects, a multitude of additional abundance peaks appear in the mass-weighted Fe abundance distribution of the gas across the LoS. These peaks consist of less metal-enriched gas and are more prominent in the systems with a higher core enhancement, since for the less enriched clusters the Fe abundance distribution remains effectively unimodal. The effect arises from large amounts of less-enriched gas that is located at larger galactocentric radii intersecting the LoS, which hence biases the mean Fe abundance low. Namely, projection effects change an average inference errors between the bvlognorm inferred metallicity and the intrinsic mass-weighted mean Fe abundance from -0.18 Solar (-28 per cent) for the central gas to 0.02 Solar (7 per cent) for the full column. The effect has a strong dependence on the mean Fe abundance of the ICM, but it is independent of halo mass, explaining the trend we see in Figure 7.
4.2.2 Brightness biases
@@
Previous studies have shown than observed ICM values are more related to emission-weighted properties than to e.g. mass-weighted properties (e.g. Biffi et al., 2017, 2018). An emission spectrum, and models to fit emission spectra, are necessarily sensitive in a weighted way towards the most luminous gas, rather than the entire mass of the ICM. We therefore quantify this bias due to brightness by comparing the mass- and emission-weighted distributions of the gas along the LoS, as well as how it influences inferences by the bvlognorm model.
@@
Starting with the thermal structure, Figure 11 shows the difference between emission and mass weighted distributions (top right). Overall, these are small, and primarily visible in the sub-keV tails. This is expected as the associated gas is located close to the cluster centre, is densest, and is therefore brighter than most other gas along the line of sight. The influence on the quality of the bvlognorm inference is similarly mild (top left). We find a shift from an average inference error of 0.16 keV (3 per cent) to -0.08 keV (-1 per cent) from the mass-weighted to the emission-weighted ICM effective temperature.
The effect is more significant for the inference of the width of the temperature distribution and yet the brightness bias does not seem to justify the differences in inference errors between the top and bottom panels of Figure 6 (not shown).
@@
Finally, in Figure 11 (bottom) we examine the effects of the brightness bias on the Fe abundance distribution (bottom right). Similar to the temperatures, by favouring the brightest gas along the line of sight, there is an enhanced contribution from the super-Solar tails, that are associated with the cluster centre, and a relatively suppressed contribution from gas in the outskirts of the cluster. As a result, the emission-weighted mean Fe abundance is biased high, resulting in an average inference error of -0.06 Solar (-18 per cent) compared to the average 0.02 Solar (7 per cent), inferred by the bvlognorm model for the mass-weighted case.
@@
In Figure 12 we present the offsets between inferred and mean, mass-weighted temperature and Fe abundance for the gas in the centres of our clusters, as a function of their central cooling time and central entropy (from Lehle et al., 2024). For both cases, we find no strong correlation between the inference error and the cool core metric, although the scatter is larger for NCC clusters.
Regarding the rejected best-fits, while a number of CC clusters fail to produce an accepted fit for all models, the rejected fits disproportionally arise from the NCC subsample. This is also evident in the spectral fit acceptance rates, as the CC cluster subsample outperforms the general population. As previously shown, constraining the temperature distribution depends on the number of counts and therefore it can be more easily accomplished the brighter the associated gas is.
Finally, we examine the associated observational inferences on the properties of the ICM, if our sample of clusters consists only of the brightest (CC) clusters. Selecting only for the systems with central cooling times below 1 Gyr (following Lehle et al., 2024), we examine their impact on temperature and Fe abundance.
In Figure 13 we show the distribution of the CC clusters versus the full sample (shown in Figures 5 and 7, for comparison). For both the temperature and Fe abundance, the average inference error increases or remains similar to that of the entire sample, respectively. For example, for the accepted bvlognorm fits, the change is from -1.14 keV (-18 per cent) to -0.69 keV (-14per cent) and from -0.15 Solar (-25 per cent) to -0.12 Solar (-23 per cent) for the mean, mass-weighted temperature and Fe abundance, respectively. However, excluding the faintest objects does remove the majority of the outliers, suggesting that the least accurate inferences occur for NCC clusters.
@@ 5 Summary and conclusions
TNG-Cluster predicts broad distributions for both temperature and Fe abundance for the gas in the centres of clusters, with a large cluster-to-cluster variation even across systems with similar total halo mass (subsection 3.1, Figure 2).
The intrinsic temperature distribution of the gas in the centre of clusters, both mass-weighted and emission-weighted, is best approximated by a Log-Normal (rather than a Normal) functional form, even if the best-fit Log-Normal description that does not fully capture the sub-keV gas tails that can be present in some, but not all, clusters (subsection 3.1, Figure 3 and Figure 4).
@@
The central ICM temperature inferred via all spectral-emission models is consistent with the true emission-weighted average along the line of sight, but systematically underestimates the mass-weighted temperature of the gas that is truly located in the clusters centre. This inference bias reads, on average (i.e. across the whole population of 352 clusters), -1.2 keV (-18 per cent), so it is arguably small, but has a strong dependence on halo mass (subsection 3.2, Figure 5). For clusters, all spectral-emission models return mean temperatures that are smaller by 2 keV, on average, than the true average one.
@@
The width of the temperature distribution inferred by the bvlognorm model, is, on average, consistent with the true mass-weighted distribution of the central ICM, but systematically underestimates the breadth of the emission-weighted distributions. In contrast, the bvgadem model overestimates the width of both mass- and emission-weighted temperature distributions (subsection 3.2, Figure 6).
@@
The inferred Fe abundance is consistent with the true mean, emission-weighted Fe abundance of the gas along the line of sight, i.e. for a central pointing with a XRISM-like FoV. The bvlognorm model is closest to the mass-weighted value. However, all models similarly underestimate the mean Fe abundance in the cluster centre by about 0.11 Solar (21 per cent) (subsection 3.3, Figure 7).
@@
The rejection of fits of the X-ray spectra correlates with the number of collected photon counts. More than half of the initially rejected bvlognorm fits do succeed with longer exposure times. However, a subsample of metal-poor systems fail to produce an accepted spectral fit, despite this longer exposure (300ks vs 100ks; subsection 4.1, Figure 9).
Projection effects are the primary reason of the offset between inferred and the actual properties of the central ICM. In contrast, the brightness bias (emission- vs. mass -weighted) has a smaller impact, at least for central pointings (subsection 4.2, Figure 10, Figure 11).
@@
When considering only strong cool-core clusters, as is the case in many observational samples, model inferences on central ICM temperature improve only marginally (with e.g. inference errors of 0.7 keV on average compared to 1.2 keV across a randomly-selected overall sample). The CC-selection effect on the inference of Fe abundance is even smaller. The main advantage of selecting for SCCs is the removal of strong outliers, i.e. of clusters for which the observational inferences of ICM properties are catastrophically wrong (subsection 4.3, Figure 13).

### v2

@@ 4.1.2 Dependence of quality of the fit on number of photons
In Figure 10 we examine this hypothesis by comparing each model inference on the temperature and Fe abundance as a function of the number of counts. It is evident from the distribution of our clusters that the rejected bvlognorm best-fits tend to have systematically lower counts than the accepted ones, but the count distributions of the two populations have a significant overlap. We note that the same behaviour holds for the accepted and rejected best-fits of the other spectral-emission models.
@@
To understand the extent that exposure time may have on the rejection of our best-fits, we repeat the X-ray spectra mocking and analysis for the rejected clusters by increasing the exposure time to 300 ks, instead of the fiducial 100 ks adopted throughout. For the case of the bvlognrom model (which returns the largest success rate), we find that 55 per cent of the clusters that have been previously rejected turn out to have a valid fit with an increased exposure (not shown). This confirms the hypothesis that the rejection of a spectral fit for the fainter clusters in our sample is due to available exposure time.
@@
However, this is not just a cluster-mass effect, as in fact there still remains a population of systems for which the best-fit is still rejected, even though their inferences on the emission-weighted mean temperature and Fe abundance are comparable with the rest of the sample. We find (although do not show) that clusters that are rejected even for a three-time longer exposure are either metal poor, with the majority of the emission-weighted mass in the ICM having an Fe abundance of 0.25 Solar, or deviate heavily from a Log-Normal distribution, with a significant amount of sub-keV gas. We note that such cases appear as almost separate populations, without any apparent correlation with halo mass (Figure 7, bottom left panel).
@@
Regarding the systems that deviate from the Log-Normal distribution, it appears that they get rejected due to their poor goodness-to-fit. It is evident that the spectral-emission model’s parametrisation is not able to fully capture their temperature distribution, leading to poor fits, despite the increased number of counts. On the other hand, the rejection of the metal poor population is due to the fact that the spectral-emission models use the emission lines also to recover the temperature distribution of the gas. Whereas the shape of the continuum is a reliable diagnostic of the ICM’s bulk temperature, for plasma with temperatures higher than 2 keV, the emission-line ratios trace the presence of cooler or hotter gas, constraining the width of the distribution. However, if clusters have low elemental abundances, they will return low numbers of counts in spectral lines irrespective of the longer exposures, and in turn the modelling will be unable to provide any meaningful constraints on the width of the temperature distribution, leading to their rejection.
@@
Overall this indicates that even the more accurate assumption of a temperature distribution is not a sufficient representation of the properties of the ICM, as predicted by TNG-Cluster, both for its thermal and more importantly its chemical structure. This highlights the limitations of current spectral-emission models and the need for more sophisticated descriptions of the intrinsic properties of the gas in order to better understand the ICM.
@@ 4.2 Effects of observational realism
Even though it might be tempting to assume that the inferred temperature and Fe abundance from a central-pointing spectrum are representative of the local conditions of the ICM, our analysis suggests otherwise. As shown e.g. in Figure 5 and Figure 7, we systematically find large inference errors on the mass-weighted state of the central ICM, but much smaller deviations from the emission-weighted properties of the projected gas along our LoS. Whereas this differentiated fit performance may not be surprising, it may have implications when searching for a physical interpretations of the data.
We now explore the individual sources of those differences and quantify their impact and systematic uncertainties. In the following sections we therefore attempt to understand and quantify the systematic uncertainties introduced by projection effects and the emission bias, by isolating the two.For the purposes of this analysis we focus again exclusively on the bvlognorm model as a reference point. However, the qualitative effects that we find are consistent across all models.
@@ 4.2.1 Projection effects
Due to inherent limitations of observations, it is not possible to differentiate emission based on the 3D position of its source. Therefore, a spectrum from a central pointing contains both the contribution from the centrally-located gas and of all the gas projected along the LoS. Previous works using simulated clusters have shown that projection effects can lead to severe misinterpretations of the intrinsic thermal and chemical properties of the ICM due to the gas along the LoS (e.g. Zhang et al., 2024; Štofanová et al., 2024).
@@
In this section we attempt to quantify the impact of the projected gas in our inference errors by comparing the mass-weighted distributions of all the cluster gas in a central pointing with a XRISM-like FoV to the mass-weighted distribution of the gas cells that actually reside in the region of interest, in this case a sphere of a 33.5 kpc radius located in the centre of each cluster.
@@
In Figure 11, we compare the temperature distribution of the gas in the centre of clusters to the total gas along the LoS observed in our mocks (top right), and their corresponding inference errors (top left). Since the temperature distribution of the gas in the clusters centre can be in most cases approximated by a unimodal distribution, we bin the clusters in percentiles based on the mode of their temperature distribution. In fact, due to the tight correlation between ICM temperature and the halo mass, the temperature bins of Figure 11 top right are comparable to mass bins. Differently than e.g. in Figure 5, here we isolate the effects of projections alone, by comparing directly mass-weighted vs. mass-weighted quantities rather than mass-weighted vs. emission weighted ones.
@@
We see that the projection of gas from larger radii along the LoS shifts the temperature distribution towards lower temperatures (dashed vs. solid curves), due the projection of cooler gas from the outskirts of the cluster. The effect is more pronounced for systems with higher ICM temperature, or equivalently higher halo mass, due to the stronger temperature gradient up to in those systems. This is consistent with the strong dependence of the temperature inference on halo mass, found in Figure 5. By comparing the difference between the inferred temperature with the bvlognorm model on the one side and the the mean, mass-weighted temperature of the central vs. LoS gas, we find much larger inference errors in the former case, with average offsets of -1.21 keV (18 per cent) and 0.15 keV (3 per cent), respectively, with a strong dependence on halo mass. Given the significant reduction in inference errors between the two, we can conclude that projection effects dominate the observational bias.
Due to the projection of cooler gas, the width of the distribution systematically increases for all systems, but the effect shows a similar positive correlation with halo mass, being more pronounced for the most massive systems. The residuals on the inferred temperature width decrease by a factor of 4, from an average of 1.05 for the central gas to 0.26 for all the gas along the LoS. This suggests that the bvlognorm model is more representative of the distribution of temperatures along the LoS rather than the level of multi-phase gas in the central regions. However, given the inference errors, the statistical significance of the residuals effectively remains, on average, within 1 .
@@
Projection effects are far more influential on the chemical structure of the ICM. In Figure 11 (bottom) we show the mass-weighted Fe abundance distributions of both central and projected gas. Similar to the temperature distributions, we bin the clusters based on the mode of the central gas, as it consists of one prominent peak. However, unlike the temperature of the ICM, TNG-Cluster shows no clear correlation between halo mass and metal enrichment. Our results indicate that, due to projection effects, a multitude of additional abundance peaks appear in the mass-weighted Fe abundance distribution of the gas across the LoS. These peaks consist of less metal-enriched gas and are more prominent in the systems with a higher core enhancement, since for the less enriched clusters the Fe abundance distribution remains effectively unimodal. The effect arises from large amounts of less-enriched gas that is located at larger galactocentric radii intersecting the LoS, which hence biases the mean Fe abundance low. Namely, projection effects change an average inference errors between the bvlognorm inferred metallicity and the intrinsic mass-weighted mean Fe abundance from -0.12 Solar (22 per cent) for the central gas to 0.08 Solar (27 per cent) for the full column. The effect has a strong dependence on the mean Fe abundance of the ICM, but it is independent of halo mass, explaining the trend we see in Figure 7.
Overall, our results indicate that systematic errors in the inferences of the temperature and metallicity of the ICM at the centre of clusters can be greatly reduced by de-projecting the observational results, irrespective of the adopted spectral-emission model. However, we note that such an approach would require observations and analysis of spectra from the full extent of the cluster, which can not be performed within the purview of a single pointing. Still, the full minimization of projection effects may still be hampered by the assumptions on e.g. the temperature distribution of the ICM along the LoS and the geometry of the cluster.
4.2.2 Effects of mass vs. emission weighting
@@
Previous studies have shown that observed ICM values are more related to emission-weighted properties than to e.g. mass-weighted properties (e.g. Biffi et al., 2017, 2018). An emission spectrum, and models to fit emission spectra, are necessarily sensitive in a weighted way towards the most luminous gas, rather than the entire mass of the ICM. As a result, physical properties of the ICM inferred by X-ray spectroscopy, such as the amount of Fe in the ICM, can be biased by the properties of the brightest gas.
In this section, we attempt to quantify how the brightest gas along the LoS contributes to the inference errors on the properties of the gas by comparing the mass- and emission-weighted distributions of the gas along the LoS, and not just that of the central sphere, as well as how it influences inferences by the bvlognorm model.
@@
Starting with the temperature, Figure 12 shows the difference between emission- and mass-weighted distributions (top right). Overall, these are small, and primarily visible in the sub-keV tails. This is expected as the associated gas is located close to the cluster centre, is densest, and is therefore brighter than most other gas along the line of sight. The influence on the quality of the bvlognorm inference is similarly mild (top left). We find a shift from an average inference error of 0.15 keV (3 per cent) to -0.08 keV (1 per cent) from the mass-weighted to the emission-weighted ICM effective temperature.
The effect on the the inference of the width of the temperature distribution due to this brightness bias is relatively small, with the residuals increasing to an average of 0.4, from 0.26 for the mass-weighted case. This further highlights the previously-found trend of the Log-Normal distribution remaining unaffected by the enhancement of the sub-keV gas. (not shown)
@@
Finally, in Figure 12 (bottom) we examine the effects of the brightness bias on the Fe abundance distribution (bottom right). Similar to the temperatures, by favouring the brightest gas along the line of sight, there is an enhanced contribution from the super-Solar tails, that are associated with the cluster centre, and a relatively suppressed contribution from gas in the outskirts of the cluster. As a result, the emission-weighted mean Fe abundance is biased high, resulting in an average inference error of 0.08 Solar (27 per cent) compared to the average 0 Solar (<1 per cent), inferred by the bvlognorm model for the mass-weighted case.
@@
In Figure 13 we present the offsets between inferred and mean, mass-weighted temperature and Fe abundance for the gas in the centres of our clusters, as a function of their central cooling time and central entropy (from Lehle et al., 2024). For all cases, we find no strong correlation between the inference error and the cool core metric, although the scatter is larger for weak cool core (WCC) and NCC clusters.
Regarding the rejected best-fits, while a number of strong CC (SCC) clusters fail to produce an accepted fit for all models, the rejected fits disproportionally arise from the NCC subsample. Namely, 70 per cent of the SCC subsample produces an accepted double bvapec and bvgadem best-fit and 80 per cent a valid bvapec and bvlognorm; these figures are much higher than the general population acceptance rates. We note that, accounting only for SCC, the bvlognorm model has the best acceptance rate, surpassing even the single temperature bvapec one. As previously shown, constraining the temperature distribution depends on the number of counts and therefore it can be more easily accomplished the brighter the associated gas is.
Finally, we examine the associated observational inferences on the properties of the ICM, if our sample of clusters consists only of SCC clusters, that are easier to detect, compared to their NCC counterparts, due to their apparent X-ray brightness (e.g. Eckert et al., 2011). Selecting only for the systems with central cooling times below 1 Gyr (following Lehle et al., 2024), we examine their impact on temperature and Fe abundance.
In Figure 14 we show the distribution of the CC clusters versus the full sample (shown in Figures 5 and 7, for comparison). For both the temperature and Fe abundance, the average inference error increases or remains similar to that of the entire sample, respectively. For example, for the accepted bvlognorm fits, the change is from -1.21 keV (18 per cent) to -0.68 keV (14 per cent) and from -0.12 Solar (22 per cent) to -0.05 Solar (9 per cent) for the mean, mass-weighted temperature and Fe abundance in the central sphere, respectively. However, excluding the faintest objects does remove the majority of the outliers, suggesting that the least accurate inferences occur for NCC clusters.
@@ 5 Summary and conclusions
The intrinsic temperature distribution of the gas in the centre of clusters, both mass-weighted and emission-weighted, can not be completely captured by either a Normal or a Log-Normal distribution. More complicated functional forms may be necessary to accurately describe the temperature distribution of the ICM in the non-negligible fraction of systems that exhibit important tails towards both cooler and super-virial temperatures (Section 3.1, Figure 3 and Figure 4).
@@
The central ICM temperature inferred via all spectral-emission models is consistent with the true emission-weighted average along the line of sight, but systematically underestimates the mass-weighted temperature of the gas that is truly located in the clusters centre. This inference bias reads, on average (i.e. across the whole population of 352 clusters), -1.2 keV (17 per cent; -0.73,+0.72 keV for the 16th-84th percentiles), so it is arguably small, but has a dependence on halo mass (Section 3.2, Figure 5). For clusters, all spectral-emission models return mean temperatures that are smaller by 1.5 keV, on average, than the true mass-weighted average one.
@@
The width of the temperature distribution inferred by the bvlognorm model, is, on average, overestimated compared to the true mass-weighted distribution of the central ICM, but within 1 of the emission-weighted distributions. In contrast, the bvgadem model overestimates the width of both mass- and emission-weighted temperature distributions to a larger extent (Section 3.2, Figure 6).
@@
The inferred Fe abundance is consistent with the true mean, emission-weighted Fe abundance of the gas along the line of sight, i.e. for a central pointing with a XRISM-like FoV. However, all models similarly underestimate the mean Fe abundance in the cluster centre by about 0.12 Solar (22 per cent), which is significantly larger than the statistical uncertainties usually associated to the measurements (Section 3.3, Figure 7).
@@
The rejection of fits of the X-ray spectra correlates with the number of collected photon counts but, interestingly, not with errors in the inference. More than half of the initially rejected bvlognorm fits do succeed with longer exposure times. However, metal-poor systems and systems with temperature distributions that deviate strongly from a Log-Normal, fail to produce an accepted spectral fit, even if we account for longer exposures (300ks vs 100ks; Section 4.1, Figure 10).
@@
When considering only strong cool-core clusters, as is the case in many observational samples, model inferences on central ICM temperature improve only marginally for the average cluster (with e.g. inference errors of 0.7 keV on average compared to 1.2 keV across a randomly-selected overall sample). The CC-selection effect on the inference of Fe abundance is even smaller. The main advantage of selecting for SCCs is the removal of strong outliers, i.e. of clusters for which the observational inferences of ICM properties are catastrophically wrong (Section 4.3, Figure 14).